# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v2

@@ 3.2 Model Weights and Updates as Linear Associative Memory Model
Figure 2: Overview of MoRA. For each new task, we freeze the ranks learned on previous tasks and introduce $r$ new ranks of updates. Our sparse self-activated mixture‐of‐ranks framework jointly considers all old and new ranks, adaptively inferring a sparse mixture weight for each rank. Panels (a,c) illustrate MoRA conceptually and (b,d) detail its computation for tasks $t$ and $t+1$, respectively.
Parameter update as modifications to the linear associative memory. Building on the associative-memory view of weights, we can likewise interpret weight updates during training as key–value insertions and modifications into the existing memory. During training, for an arbitrary update $\Delta\mathbf{W}$ on a given $\mathbf{W}_{0}$ as the starting point, the weight update process can be represented as
$\mathbf{W}_{\text{new}}=\mathbf{W}_{0}+\Delta\mathbf{W},$ (3)
where $\mathbf{W}_{0}$ can be the pre-trained model weights. We can also formulate the weight updates $\Delta\mathbf{W}$ with the key-value memory model $\{\Delta\mathbf{K},\Delta\mathbf{V}\}$ as an update to the original memory model. With the operation on an input $\mathbf{x}$ via
$\Delta\mathbf{W}\mathbf{x}=\Delta\mathbf{V}\Delta\mathbf{K}\mathbf{x},$ (4)
the updates on weights can be seen as a retrievable key-value system for the updating of $\mathbf{W}\mathbf{x}$.
LoRA as low-rank key–value memory updates. LoRA (Hu et al., 2021) performs parameter‐efficient fine‐tuning by adding a rank‐$r$ update $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ to the pre‐trained weight matrix $\mathbf{W}_{0}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$. With the model update based on LoRA, given an input $\mathbf{x}$, the computation of updated weight becomes
$\mathbf{y}_{\text{new}}=\mathbf{W}_{\text{new}}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}.$ (5)
Decomposing the low-rank matrices into a collection of its $r$ rank-one components gives
$\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{:,i}(\mathbf{A}_{i,:}\mathbf{x}),$ (6)
with $\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1,:}^{\top}&\cdots&\mathbf{A}_{r,:}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\text{in}}}$ where $\mathbf{A}_{i,:}$ is the $i$-th row of $\mathbf{A}$, and $\mathbf{B}=\begin{bmatrix}\mathbf{B}_{:,1}&\cdots&\mathbf{B}_{:,r}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times r}$ where $\mathbf{B}_{:,i}$ is the $i$-th column of $\mathbf{B}$.
Considering the connection between the formulation of LoRA Eq. (1) and associative‐memory view (Anderson, 1972; Meng et al., 2022) of weight update Eq. (4), we can intuitively interpret LoRA as obtaining a small key–value store in a low‐rank subspace, where the rows of $\mathbf{A}$ serve as keys and the columns of $\mathbf{B}$ serve as their corresponding values. After training the LoRA module on a base weight matrix $\mathbf{W}_{0}$, the modified forward pass in Eq. (5) with Eq. (6) can be seen as: first, the original pre‐trained computation $\mathbf{W}_{0}\,\mathbf{x}$ (Eq. (2)) is performed; then, each learned key $\mathbf{A}_{i,:}$ measures its relevance to the input $\mathbf{x}$ via the product $\mathbf{A}_{i,:}\mathbf{x}$, and each learned value $\mathbf{B}_{:,i}$ is scaled by this relevance score and aggregated, in a same way as shown in Eq. (2). In this way, LoRA effectively inserts or updates a small set of key–value memories that augment the pre‐trained model’s capabilities.
3.3 Mixture of Ranks Adaptation
In the key–value associative‐memory view of LoRA (Eq. (5) and Eq. (6)), a standard LoRA or its variants (Hu et al., 2021; Meng et al., 2024; Zhang et al., 2024a) apply a dense update: every rank’s subspace contributes to $\Delta\mathbf{W}\,\mathbf{x}$ for any input $\mathbf{x}$. Even in MoE-LoRA, where each LoRA module is treated as a separate expert, all key–value pairs within a LoRA module are typically used densely as a whole (either fully activated or not), which creates redundancy across experts and underutilizes the expressive potential of the key vectors $\mathbf{A}$ to indicate input relevance.
Rather than by default activating all key–value memories for every input, many of which may encode patterns irrelevant to the current token, we introduce a gated, sparse mixture-of-ranks update that selectively activates only the most relevant rank-one key–value pairs. Concretely, we associate each rank-one key–value pair with a scalar weight ${\textnormal{w}}_{i}\in\mathbb{R}$. For an input token $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, with $r_{t}$ total ranks, the update becomes
@@ 3.3 Mixture of Ranks Adaptation
where ${\textnormal{w}}_{i}$ modulates whether and how strongly each rank $i$ contributes to the updated weight matrix.
3.4 Self-Activated Adaptive Sparse Mixture of Ranks
A key challenge for MoE-style CL is routing ambiguity, leading to catastrophic forgetting. Conventional routers assign examples to coarse-grained experts, but adding new experts often makes routing suboptimal, causing over-activation or underuse, i.e., the three challenges discussed in Sec. 1. To avoid these pitfalls, we treat each rank-one LoRA (or rank) as an expert based on the rank-level mixture model (Eq. (7)), enabling sparse activation of fine-grained experts for specific tasks. Thus, CL with incrementally added rank-one experts becomes a process of accumulating atomic key–value knowledge that can be reused in future tasks. Unlike conventional MoE-LoRA (Yang et al., 2024a; Wang et al., 2024; Yu et al., 2024), we use the rank-one $\mathbf{A}_{i}$’s (keys) as routers to their corresponding $\mathbf{B}_{i}$’s (values) through self-activation, with each $\mathbf{A}_{i}$$\mathbf{B}_{i}$ pair functioning as a router-expert pair.
Mixture of self-activated rank-one experts. Rather than relying on an external router, our framework derives mixture weights for each rank directly from that rank’s activation. Recall from Eq. (6) that each rank-one update is parameterized by a key–value memory pair $(\mathbf{A}_{i,:},\mathbf{B}_{:,i})_{i=1}^{r_{t}}$; the key $\mathbf{A}_{i,:}$ therefore provides a natural, input-dependent relevance signal. Specifically, given an input token $x\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, and treating $\mathbf{A}$ as the set of keys, we compute a raw score for all $r_{t}$ ranks accumulated up to task $t$:
${\textnormal{s}}_{i}\;=\;\frac{\mathbf{A}_{i}x}{\sqrt{\sum_{j=1}^{r_{t}}\|\mathbf{A}_{j}x\|_{2}^{2}}}\quad\in\mathbb{R},$ (8)
where the $\ell_{2}$-normalization across all $r_{t}$ ranks ensures numerical stability and compensates for activation-scale variations between layers and pre-trained weights. Using raw scores alone matches or surpasses a separately learned router (with 1.5 times parameters) without self-activation (Table 4), showing that each rank can infer its own relevance from activation.
Self-activated rank selection. To ensure a stable and gated, sparse activation pattern, we impose a fixed sparse rank‐activation budget via top-k masking on the $\ell_{2}$–normalized scores ${\textnormal{s}}_{i}:=\mathrm{TopK}({\textnormal{s}},\,k),$ where
@@ 3.4 Self-Activated Adaptive Sparse Mixture of Ranks
ensuring that at most $k$ out of the total $r_{t}$ ranks are eligible to be activated for task $t$. This encourages rank specialization: only a small set of the most relevant ranks are emphasized and trained to capture each kind of input-specific dynamics, and it prevents tiny, noisy activations from affecting gradients and learning dynamics.
Sharpness control on rank activations. To further encourage rank specialization and concentrate the update on the most relevant ranks, we temperature-scale and normalize the scores:
${\textnormal{w}}_{i}\;=\;\mathrm{softmax}\!\bigl(\tfrac{s}{\tau_{\text{MoRA}}}\bigr)_{i},$ (10)
where the softmax is taken over the vector ${\textnormal{s}}=[{\textnormal{s}}_{1},\dots,{\textnormal{s}}_{r_{t}}]$ and $\tau_{\text{MoRA}}$ controls the distribution’s sharpness. A lower $\tau_{\text{MoRA}}$ concentrates mass on a few ranks (specialists), sharpens the distribution and routes larger learning signals to top ranks, which accelerates specialization and reduces noisy updates to low-relevance ranks, while a higher $\tau_{\text{MoRA}}$ spreads mass (shared representations).
Threshold-based test-time rank selection. To further suppress low-signal ranks at inference, we optionally apply a threshold $\delta$ to the $\ell_{2}$–normalized raw scores and zero out gates below it. We do not hard-threshold during training such that gradients can still flow to near-threshold ranks and allow exploration and adaptation; thresholding only at test time reduces noise and runtime cost:
@@
This yields a sparse, input-dependent mixture of only the most significant rank-one experts.
Our fine-grained mixture framework offers two key advantages: (1) It reduces redundancy and enables shared-knowledge reuse: ranks that capture common, cross-task features (such as generic semantics) can be reactivated across multiple tasks, preventing those patterns from being relearned redundantly or interfered; (2) It minimizes routing ambiguity and prevents forgetting: each rank can specialize on more nuanced patterns, yielding an optimized, rank‐wise mixture that integrates new knowledge without disrupting existing representations or hindering future updates.
Incremental mixture of ranks. Unlike methods that add a fixed number of LoRA experts per task (Yu et al., 2024), MoRA incrementally adds $r$ new ranks for each task, freezes all prior ranks, and adaptively selects a sparse subset of the most relevant ranks during both training and inference. This design seamlessly incorporates new task-specific knowledge into freshly initialized ranks while preserving earlier ranks unchanged to prevent forgetting. Moreover, it facilitates common knowledge reuse by allowing ranks that encode common patterns to be activated in later tasks.
Training Objectives. In our experiments, we only use the model’s standard training objective, without any extra regularization or load‐balancing constraints. In practice, we empirically observe that the conventional load-balancing loss widely employed in the Mixture-of-Experts literature (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024) is not necessary, thanks to our fine-grained self-activated gating design. Imposing a strict load-balancing constraint might hinder the learning of the key vectors (i.e., matrix $\mathbf{A}$), since each rank should remain free to capture its own characteristic input distribution. Fig. 3 illustrates how each rank learns distinct characteristics.
@@
Method
Aircraft Caltech DTD EuroSAT Flowers Food MNIST OxPet Cars SUN397 Average
@@
   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF (Li & Hoiem, 2017) 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT (Wortsman et al., 2022) 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL (Rebuffi et al., 2017) 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL (Zheng et al., 2023) 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
   MoE-Adapter (Yu et al., 2024) 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
   RAIL-Primal (Xu et al., 2024) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   CoDyRA (Lu et al., 2024) 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
   MoRA 74.5 38.1 46.9 65.3 82.9 45.8 88.2 65.1 62.9 63.3
@@
   LwF (Li & Hoiem, 2017) 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT (Wortsman et al., 2022) 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL (Rebuffi et al., 2017) 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL (Zheng et al., 2023) 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
   MoE-Adapter (Yu et al., 2024) 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
   RAIL-Primal (Xu et al., 2024) 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
   CoDyRA (Lu et al., 2024) 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
   MoRA 44.1 81.6 64.6 79.6 83.9 84.4 66.5 89.7 68.4 64.1 72.7
@@
   LwF (Li & Hoiem, 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT (Wortsman et al., 2022) 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL (Rebuffi et al., 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL (Zheng et al., 2023) 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
   MoE-Adapter (Yu et al., 2024) 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
   RAIL-Primal (Xu et al., 2024) 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
   CoDyRA (Lu et al., 2024) 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
   MoRA 37.7 81.5 70.7 92.4 95.0 86.0 97.6 92.6 81.0 74.7 80.9

### v3

@@ 3.1 Preliminaries
Figure 2: Overview of MoRAM. For each new task, we freeze the ranks learned on previous tasks and introduce $r$ new ranks of updates. Our sparse self-activated mixture‐of‐ranks framework jointly considers all old and new ranks, adaptively inferring a sparse mixture weight for each rank. Panels (a,c) illustrate MoRAM conceptually and (b,d) detail its computation for tasks $t$ and $t+1$, respectively.
3.2 Weights as Linear Associative Memory
To resolve the limitations of coarse granularity and routing ambiguity inherent in dense rank-$r$ updates, we depart from the conventional view of LoRA as monolithic blocks, instead reconceptualizing both the weight matrix and its updates as a Linear Associative Memory.
Definition 1 (Weight Matrix as Linear Associative Memory).
Consider a weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ of rank $m$. Through the lens of linear associative memory [37, 5], a weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ of rank $m$ acts as a retrieval system composed of $m$ atomic key-value pairs $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{m}$, where $\mathbf{k}_{i}\in\mathbb{R}^{d_{\text{in}}}$ and $\mathbf{v}_{i}\in\mathbb{R}^{d_{\text{out}}}$ such that $\mathbf{W}\approx\sum_{i=1}^{m}\mathbf{v}_{i}\mathbf{k}_{i}^{\top}$. For an input hidden state for a token $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$, the matrix-vector product effectively performs a content-addressable read operation:
$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{v}_{i}(\mathbf{k}_{i}^{\top}\mathbf{x}),$ (2)
where the inner product $(\mathbf{k}_{i}^{\top}\mathbf{x})$ computes the relevance (or activation strength) of the $i$-th memory slot to the input, which weights the retrieval of the value vector $\mathbf{v}_{i}$.
Remark 2 (Key-Value Memory of Weight Matrix).
It is crucial to distinguish Definition 1 from the key-value mechanism in Self-Attention [73]. In attention, keys and values are dynamic projections of the input sequence generated at runtime. In contrast, under the linear associative memory view, the keys $\mathbf{k}_{i}$ and values $\mathbf{v}_{i}$ are static parameters intrinsic to the weight matrix, representing knowledge patterns acquired during pre-training.
3.3 Proposed MoRAM
3.3.1 Mixture of Rank-1 Memory Experts
Fine-grained rank-1 memory augmentation. Grounded by the associative memory view (Definition 1), we reconceptualize the fine-tuning process. Instead of formulating the update $\Delta\mathbf{W}$ as a single-unit rank-$r$ matrix, we treat it as a collection of retrievable rank-1 memory augmentations. Formally, we define the update as the aggregation of $r$ atomic rank-1 key-value pair updates:
$\Delta\mathbf{W}\mathbf{x}=\sum\nolimits_{i=1}^{r}\underbrace{\mathbf{B}_{:,i}}_{\text{Value }\mathbf{v}_{i}}(\underbrace{\mathbf{A}_{i,:}}_{\text{Key }\mathbf{k}_{i}^{\top}}\mathbf{x}),$ (3)
where the row vector $\mathbf{A}_{i,:}$ acts as the Key determining the relevance of the atom to the input, and the column vector $\mathbf{B}_{:,i}$ acts as the Value storing the retrieved knowledge. Crucially, this shifts the paradigm from matrix adaptation to memory expansion: the update is no longer a rigid block, but a flexible set of fine-grained atomic memories.
Despite this granular potential, standard LoRA and its variants [33, 56, 98] apply updates densely: every rank contributes to every input. Even in MoE-LoRA, while experts are separated at rank-$r$ adapter level, the constituent rank-1 memories within each expert remain entangled: (1) Interference: Irrelevant knowledge subspaces within a chosen expert are forced to activate for mismatched inputs. (2) Routing collapse: It overlooks the intrinsic capacity of the key vectors $\mathbf{A}$ to act as content retrieval keys, instead relying on indiscriminate dense activation or redundant external routers.
MoRAM formulation. To address this, we propose MoRAM (Mixture of Rank-1 Associative Memory). We dismantle the rigid adapter structure, redefining the model’s adaptation parameters as a dynamic collection of atomic memory experts $\mathcal{M}_{t}=\{(\mathbf{B}_{:,i},\mathbf{A}_{i,:})\}_{i=1}^{r_{t}}$, where $r_{t}$ denotes the total accumulated rank-1 pairs avaliable at task $t$. For an input $\mathbf{x}$ during task $t$ (with accumulated memory atoms $r_{t}$), the effective update is defined as a sparse, input-dependent mixture:
@@ 3.3.1 Mixture of Rank-1 Memory Experts
where ${\textnormal{w}}_{i}\in\mathbb{R}$ represents the computed retrieval confidence of the $i$-th memory atom on each input token hidden state $\mathbf{x}$}. This formulation allows the model to freeze specific “atomic” memories (preserving old tasks) while inserting new ones on demand of input, or to jointly activate a combination of old and new memory slots to handle shared concepts.
3.3.2 Self-Activation for MoRAM Routing
Relying on the key-value form of the rank-1 adapter, we propose a self-activation mechanism for routing the memory experts. We derive the mixing weights directly from the intrinsic activation of each atomic memory, leveraging the key vectors $\mathbf{A}_{i,:}$ for memory retrieval. Unlike standard MoE approaches that rely on additional routing networks, our self-activation routing performs routing as a content-addressable retrieval, reducing the possibility of forgetting caused by an additional router.
Self-activated relevance scoring. By eliminating the external router, MoRAM avoids the optimization instability and forgetting associated with auxiliary networks. Instead, we derive mixing weights directly from the intrinsic alignment between the input and the memory keys. Formally, for an input token $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ and the set of accumulated memory atoms up to task $t$, we compute the raw relevance score ${\textnormal{s}}_{i}$ for the $i$-th atom as:
${\textnormal{s}}_{i}\;=\;\frac{\mathbf{A}_{i,:}\mathbf{x}}{\sqrt{\sum_{j=1}^{r_{t}}(\mathbf{A}_{j,:}\mathbf{x})^{2}}},$ (5)
where the numerator represents the relevance to the memory key $\mathbf{A}_{i,:}$, the denominator performs an $\ell_{2}$-normalization across the entire memory ensemble for numerical stability. Empirically, we find that this intrinsic scoring mechanism matches or exceeds the performance of external routers (see Table 4), confirming that the memory keys alone contain sufficient information to determine their own utility.
3.3.3 Sparse Expert Routing and Mixture
While Eq. (5) measures relevance, naive dense activation leads to low specialization and induces interference and computational overhead. We therefore employ sparse routing to enhance fine-grained specialization of the fine-grained experts and routers.
Sparse rank selection. To prevent interference between tasks and ensure a budgeted computational cost, we enforce a sparsity constraint via top-k masking. Given the raw relevance scores s, we retain only the $k$ entries with the highest activation and mask the rest:
@@ 3.3.3 Sparse Expert Routing and Mixture
This masking ensures that for any given input, at most $k$ out of the total $r_{t}$ accumulated memory atoms are eligible for activation. This encourages rank specialization: only a small set of the most relevant ranks are emphasized and trained to capture each kind of input-specific dynamics, and it prevents tiny, noisy activations from spuriously triggering unrelated memory atoms (e.g., those frozen from prior tasks).
Sharpness enhancement. To further encourage rank specialization and concentrate the update on the most relevant ranks, we apply temperature-scaled softmax normalization to enhance the sharpness of the mixture weights:
${\textnormal{w}}_{i}\;=\;\text{softmax}\Bigl(\frac{\mathrm{TopK}({\textnormal{s}},k)}{\tau_{\text{MoRAM}}}\Bigr)_{i},$ (7)
where $\tau_{\text{MoRAM}}$ is a scalar temperature hyperparameter. In the forward pass, a lower $\tau_{\text{MoRAM}}$ acts as a contrast enhancer, concentrating probability mass on the definitive specialist ranks. In the backward pass, it functions as a gradient router: by sharpening the distribution, stronger learning signals are directed exclusively to the winning ranks. This effectively isolates them from irrelevant updates and accelerates specialization, ensuring that memory atoms only evolve when they are truly relevant to the current data distribution.
Threshold-based expert selection. To maximize retrieval precision during inference, we apply an additional threshold-based expert selection, via a relevance threshold $\delta$ to the normalized scores, enabling adaptively pruning memory atoms with weak activation signals. Since the number of useful experts may differ depending on layers and input, at inference time, we apply this threshold-based filter to further adaptively select experts from the top-k experts, while reducing both computational overhead and noise:
@@
This operation yields a highly sparse, input-dependent set comprising only the most significant memory experts.
3.3.4 Continual Learning with MoRAM
Incremental memory expansion. Unlike methods that add fixed-size adapter blocks [96], MoRAM expands memory “little by little.” For each new task, we introduce a set of $r$ new atomic rank-1 pairs (Keys $\mathbf{A}$ and Values $\mathbf{B}$), freeze all prior atoms, and let the self-activation mechanism jointly route across the union of all old and new memories. This seamlessly integrates new knowledge while preserving the reusability of prior tasks. To enhance efficiency and enable sub-linear expansion rate, MoRAM supports low-utility atoms to be pruned after training with minimal performance degradation (more in Appendix A.10).
Training objectives. In our experiments, we only use the model’s standard training objective, without any extra regularization or load‐balancing constraints. Benefiting from the fine-grained modeling with rank-1 experts, the experts tend to be special and the routing tends to be balanced and stable naturally. Even without additional regularization, the specialized experts and router reduce router degradation and forgetting, while we believe further regularization might be more helpful. In practice, we empirically observe that conventional load-balancing regularization [68, 40, 24, 15] for MoE may not be necessary for the experimental setup. And imposing regularization on the self-activation routing improperly tends to hinder the learning of the key vectors, since the key vectors for routing are also crucial for information representation.
@@ 3.3.4 Continual Learning with MoRAM
    Method Aircraft Caltech DTD EuroSAT Flowers Food MNIST OxPet Cars SUN397 Average
@@
   Zero-shot [63] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF [44] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT [86] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL [66] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL [105] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
   MoE-Adapter [96] 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
   RAIL-Primal [91] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   CoDyRA [50] 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
   MoRAM 74.5 38.1 46.9 65.3 82.9 45.8 88.2 65.1 62.9 63.3
@@
   LwF [44] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT [86] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL [66] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL [105] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
   MoE-Adapter [96] 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
   RAIL-Primal [91] 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
   CoDyRA [50] 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
   MoRAM 44.1 81.6 64.6 79.6 83.9 84.4 66.5 89.7 68.4 64.1 72.7
@@
   LwF [44] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT [86] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL [66] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL [105] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
   MoE-Adapter [96] 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
   RAIL-Primal [91] 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
   CoDyRA [50] 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
   MoRAM 37.7 81.5 70.7 92.4 95.0 86.0 97.6 92.6 81.0 74.7 80.9